# Early Directional Convergence in Deep Homogeneous Neural Networks for Small Initializations

### v1

@@ A.3 Gronwall’s Inequality
$\frac{\mathbf{w}_{z}(\overline{T}_{\epsilon})^{\top}{\mathbf{u}}^{*}}{\|\mathbf{w}_{z}(\overline{T}_{\epsilon})\|_{2}}$ $=\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}}^{*}+\zeta^{\top}{\mathbf{u}}^{*}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}$
$=\left(\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}}^{*}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\right)\frac{\|\mathbf{u}(T_{\epsilon})\|_{2}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}+\frac{\zeta^{\top}{\mathbf{u}}^{*}}{\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}}$
where in the first inequality we used (83) along with the fact that $\|\zeta\|_{2}\leq\epsilon$ and $\|\mathbf{u}(T_{\epsilon})+\zeta\|_{2}\geq\eta/2.$ The second inequality follows since $\|\zeta\|_{2}\leq\epsilon$ and $\|\mathbf{u}(T_{\epsilon})\|_{2}\geq\eta.$
C.2 Proof of Lemma 5.4
Proof.
We begin by proving that $\overline{\mathbf{w}}_{n}$ is a saddle point of
$\mathcal{L}\left(\mathbf{w}_{n},\mathbf{0}\right)=\frac{1}{2}\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})-\mathbf{y}\right\|^{2}.$ (85)
Note that $\nabla_{\mathbf{w}_{n}}\mathcal{L}\left({\mathbf{w}}_{n},\mathbf{0}\right)=\mathcal{J}_{n}(\mathbf{X};{\mathbf{w}}_{n})^{\top}\left(\mathcal{H}_{n}(\mathbf{X};{\mathbf{w}}_{n})-\mathbf{y}\right).$ Since $\{\overline{\mathbf{w}}_{n},\overline{\mathbf{w}}_{z}\}$ is a saddle point of
$\mathcal{L}\left(\mathbf{w}_{n},\mathbf{w}_{z}\right)=\frac{1}{2}\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z})-\mathbf{y}\right\|^{2},$ (86)
we have
$\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\end{bmatrix}=\begin{bmatrix}\mathcal{J}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})^{\top}\overline{\mathbf{y}}\\ \mathcal{J}_{z}(\mathbf{X};\overline{\mathbf{w}}_{z})^{\top}\overline{\mathbf{y}}\end{bmatrix},$ (91)
where $\overline{\mathbf{y}}=\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})+\mathcal{H}_{z}(\mathbf{X};\overline{\mathbf{w}}_{z})$. Since $\overline{\mathbf{w}}_{z}=\mathbf{0}$ implies $\mathcal{H}_{z}(\mathbf{X};\overline{\mathbf{w}}_{z})=0$, we have
@@
$\mathcal{J}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})^{\top}\left(\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})\right)=\mathbf{0}\implies\nabla_{\mathbf{w}_{n}}\mathcal{L}\left(\overline{\mathbf{w}}_{n},\mathbf{0}\right)=\mathbf{0},$
proving our first result. Next, suppose there exists $\gamma>0$ such that
$\left\langle\mathbf{w}_{n}-\overline{\mathbf{w}}_{n},\nabla_{\mathbf{w}_{n}}\mathcal{L}\left({\mathbf{w}}_{n},\mathbf{0}\right)\right\rangle\geq 0,\text{ if }\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma.$ (92)
Choose any $\mathbf{w}_{n}$ such that $\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma$. By the mean value theorem we have
$\mathcal{L}\left(\overline{\mathbf{w}}_{n},\mathbf{0}\right)=\mathcal{L}\left({\mathbf{w}}_{n},\mathbf{0}\right)+\left\langle\nabla_{\mathbf{w}_{n}}\mathcal{L}\left(\tilde{\mathbf{w}}_{n},\mathbf{0}\right),\overline{\mathbf{w}}_{n}-{\mathbf{w}}_{n}\right\rangle,$
where $\tilde{\mathbf{w}}_{n}=(1-t)\overline{\mathbf{w}}_{n}+t{\mathbf{w}}_{n}$, for some $t\in(0,1)$. Note that ${\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}=\left(\tilde{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\right)/t$. Thus,
$\mathcal{L}(\overline{\mathbf{w}}_{n},\mathbf{0})=\mathcal{L}(\mathbf{w}_{n},\mathbf{0})-\frac{1}{t}\left\langle\nabla_{\mathbf{w}_{n}}\mathcal{L}(\tilde{\mathbf{w}}_{n},\mathbf{0}),\tilde{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\right\rangle\leq\mathcal{L}(\mathbf{w}_{n},\mathbf{0}),$
where the inequality follows from (92), since $\|\tilde{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\|_{2}=t\|{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma$. Thus, $\overline{\mathbf{w}}_{n}$ is a local minimum. We next move towards proving the final part. Since $\overline{\mathbf{w}}_{n}$ is a local minimum of $\mathcal{L}({\mathbf{w}}_{n},\mathbf{0})$, there exists some $\gamma_{1}>0$ such that
$\mathcal{L}(\overline{\mathbf{w}}_{n},\mathbf{0})\leq\mathcal{L}({\mathbf{w}}_{n},\mathbf{0}),\text{ if }\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma_{1}.$
Further, since $\nabla^{2}_{\mathbf{w}_{n}}\mathcal{L}(\overline{\mathbf{w}}_{n},\mathbf{0})\succ\mathbf{0}$ and $\nabla^{2}_{\mathbf{w}_{n}}\mathcal{L}({\mathbf{w}}_{n},\mathbf{0})$ is continuous, there exists some $\gamma_{2}$ such that $\nabla^{2}_{\mathbf{w}_{n}}\mathcal{L}({\mathbf{w}}_{n},\mathbf{0})\succeq\mathbf{0},$ if $\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma_{2}$. Let $\gamma=\min(\gamma_{1},\gamma_{2})$, and choose any $\mathbf{w}_{n}$ such that $\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma$. By Taylor’s theorem we have
$\mathcal{L}(\overline{\mathbf{w}}_{n},\mathbf{0})=\mathcal{L}({\mathbf{w}}_{n},\mathbf{0})+\left\langle\nabla_{\mathbf{w}_{n}}\mathcal{L}({\mathbf{w}}_{n},\mathbf{0}),\overline{\mathbf{w}}_{n}-{\mathbf{w}}_{n}\right\rangle+\frac{1}{2}\left(\overline{\mathbf{w}}_{n}-{\mathbf{w}}_{n}\right)^{\top}\nabla^{2}_{\mathbf{w}_{n}}\mathcal{L}(\tilde{\mathbf{w}}_{n},\mathbf{0})\left(\overline{\mathbf{w}}_{n}-{\mathbf{w}}_{n}\right),$
where $\tilde{\mathbf{w}}_{n}=(1-t)\overline{\mathbf{w}}_{n}+t{\mathbf{w}}_{n}$, for some $t\in(0,1)$. Since $\|\tilde{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\|_{2}=t\|{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma$, we have $\nabla^{2}\mathcal{L}(\tilde{\mathbf{w}}_{n},\mathbf{0})\succeq\mathbf{0}$. Further, $\mathcal{L}(\overline{\mathbf{w}}_{n},\mathbf{0})\leq\mathcal{L}({\mathbf{w}}_{n},\mathbf{0})$, since $\|{\mathbf{w}}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma\leq\gamma_{1}$. Thus,

### v2

@@ C.3 Proof of Theorem 4
${\mathbf{Z}}_{l}$ $\propto\text{diag}(\sigma^{\prime}(\overline{{\mathbf{W}}}_{l}{\mathbf{a}}_{l-1}^{p})){\mathbf{b}}_{l+1}({\mathbf{a}}_{l-1}^{p})^{\top}\propto\text{diag}({\mathbf{a}}_{l}^{p-1}){\mathbf{b}}_{l+1}({\mathbf{a}}_{l-1}^{p})^{\top}.$
Now, if $p$ is odd, then $p-1$ is even. Since the entries of ${\mathbf{a}}_{l-1}$ have identical absolute values, $({\mathbf{a}}_{l-1}^{p-1})$ have identical entries, and thus, ${\mathbf{a}}_{l-1}^{p}=\text{diag}({\mathbf{a}}_{l-1}^{p-1}){\mathbf{a}}_{l-1}\propto{\mathbf{a}}_{l-1}.$ Also, since ${\mathbf{b}}_{l}=q_{l-1}{\mathbf{a}}_{l-1}/p^{1/4}$, we have ${\mathbf{a}}_{l-1}^{p}\propto{\mathbf{b}}_{l}$. Furthermore, since ${\mathbf{b}}_{l+1}=q_{l}{\mathbf{a}}_{l}/p^{1/4}$ and the entries of ${\mathbf{a}}_{l}$ have identical absolute values, $\text{diag}({\mathbf{a}}_{l}^{p-1}){\mathbf{b}}_{l+1}\propto{\mathbf{b}}_{l+1}\propto{\mathbf{a}}_{l}$. Hence ${\mathbf{Z}}_{l}\propto{\mathbf{a}}_{l}{\mathbf{b}}_{l}^{\top}\propto\overline{{\mathbf{W}}}_{l}.$ Next, if $p$ is even, then ${\mathbf{a}}_{l-1}^{p}=|{\mathbf{a}}_{l-1}|^{p}$. Since the entries of ${\mathbf{a}}_{l-1}$ have identical absolute values, $|{\mathbf{a}}_{l-1}|^{p-1}$ have identical entries, and thus, $|{\mathbf{a}}_{l-1}|^{p}\propto|{\mathbf{a}}_{l-1}|.$ Also, since ${\mathbf{b}}_{l}=q_{l-1}|{\mathbf{a}}_{l-1}|/p^{1/4}$, we have ${\mathbf{a}}_{l-1}^{p}\propto{\mathbf{b}}_{l}$. Furthermore, since ${\mathbf{b}}_{l+1}=q_{l}|{\mathbf{a}}_{l}|/p^{1/4}$, we have $\text{diag}({\mathbf{a}}_{l}^{p-1}){\mathbf{b}}_{l+1}\propto{\mathbf{a}}_{l}$, which implies ${\mathbf{Z}}_{l}\propto{\mathbf{a}}_{l}{\mathbf{b}}_{l}^{\top}\propto\overline{{\mathbf{W}}}_{l}.$
Case 2 $(\alpha\neq 1):$ We begin by proving the “only if” part. Since $\sigma(\cdot)$ is $p-$positively homogeneous, therefore,
$\sigma(c{\mathbf{p}})=c^{p}\sigma({\mathbf{p}}),\sigma^{\prime}({\mathbf{q}})=p{\mathbf{q}}^{p-1}\text{ and }\sigma(d{\mathbf{q}})={\mathbf{q}}^{p}\sigma(d),$ (111)
where $c>0$, ${\mathbf{p}}$ is a vector, and $d\in{\mathbb{R}}$, ${\mathbf{q}}$ is a vector with non-negative entries. Also, we use $(\sigma^{q})^{\prime}(\cdot)$ to denote the derivative of $\sigma^{q}(\cdot)$, where $q\in{\mathbb{N}}.$
From Lemma 5, for all $l\in[L-2]$,
$|{\mathbf{a}}_{l}|={p^{1/4}}|{\mathbf{b}}_{l+1}|,\text{ and }|{\mathbf{a}}_{L-1}|={(p\hat{p})^{1/4}}|\overline{{\mathbf{w}}}|.$ (112)
Now, since the entries of $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ are non-negative, from eq. 112, we have ${\mathbf{b}}_{l}={\mathbf{a}}_{l-1}/p^{1/4}$, for $12\leq l\leq L-$ 1, and $|\overline{{\mathbf{w}}}|={\mathbf{a}}_{L-1}/(p\hat{p})^{1/4}$.
We next show that, for $l\in[L-1]$, each non-zero entry of ${\mathbf{a}}_{l}$ has identical values. Let $\phi_{i}^{l}=\sigma(\overline{{\mathbf{W}}}_{l}\cdots\sigma(\overline{{\mathbf{W}}}_{1}{\mathbf{x}}_{i})\cdots)$, for $l\in[L-1]$, for $11\leq l\leq L-$ 1, then, using eq. 111,
$\phi_{i}^{l}=\sigma(\overline{{\mathbf{W}}}_{l}\cdots\sigma(\overline{{\mathbf{W}}}_{1}{\mathbf{x}}_{i})\cdots)$ $=\sigma({\mathbf{a}}_{l}{\mathbf{b}}_{l}^{\top}\sigma({\mathbf{a}}_{l-1}{\mathbf{b}}_{l-1}^{\top}\cdots\sigma({\mathbf{a}}_{2}{\mathbf{b}}_{2}^{\top}\sigma({\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top}{\mathbf{x}}_{i}))\cdots)$
$={\mathbf{a}}_{l}^{p}\sigma({\mathbf{b}}_{l}^{\top}\sigma({\mathbf{a}}_{l-1}{\mathbf{b}}_{l-1}^{\top}\cdots\sigma({\mathbf{a}}_{2}{\mathbf{b}}_{2}^{\top}\sigma({\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top}{\mathbf{x}}_{i}))\cdots)={\mathbf{a}}_{l}^{p}c_{i}^{l},$
where $c_{i}^{l}:=\sigma({\mathbf{b}}_{l}^{\top}\sigma({\mathbf{a}}_{l-1}{\mathbf{b}}_{l-1}^{\top}\cdots\sigma({\mathbf{a}}_{2}{\mathbf{b}}_{2}^{\top}\sigma({\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top}{\mathbf{x}}_{i}))\cdots)$ is a scalar. Since $(\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})$ is a non-zero KKT point of eq. 10, from Lemma 16 we have, for $12\leq l\leq L-$ 1,
@@
$\mathbf{0}=\lambda\overline{{\mathbf{W}}}_{l}+\nabla_{{\mathbf{W}}_{l}}\left({\mathbf{y}}^{\top}{\mathcal{H}}({\mathbf{X}};\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})\right)$ $=\lambda{\mathbf{a}}_{l}{\mathbf{b}}_{l}^{\top}+\sum_{i=1}^{n}{\mathbf{e}}_{i}^{l}\left(\phi_{i}^{l-1}\right)^{\top}$
$=\lambda{\mathbf{a}}_{l}{\mathbf{b}}_{l}^{\top}+\sum_{i=1}^{n}{\mathbf{e}}_{i}^{l}c_{i}^{l-1}\left({\mathbf{a}}_{l-1}^{p}\right)^{\top}.$ (113)
Multiplying the above equation by ${\mathbf{a}}_{l}^{\top}$ from the left gives us
$0=\lambda\|{\mathbf{a}}_{l}\|_{2}^{2}{\mathbf{b}}_{l}^{\top}+\left(\sum_{i=1}^{n}{\mathbf{a}}_{l}^{\top}{\mathbf{e}}_{i}^{l}c_{i}^{l-1}\right)\left({\mathbf{a}}_{l-1}^{p}\right)^{\top}.$ (114)
The above equation implies that ${\mathbf{b}}_{l}$ is parallel to ${\mathbf{a}}_{l-1}^{p}$. Hence, from eq. 112, ${\mathbf{a}}_{l-1}$ is parallel to ${\mathbf{a}}_{l-1}^{p}$. Now, since $p\geq 2$, and ${\mathbf{a}}_{l}$ has non-negative entries, we get that, for $21\leq l\leq L-$ 2, the non-zero entries of ${\mathbf{a}}_{l}$ are identical.
From the KKT condition for $l=L$, we have
$\mathbf{0}=\lambda\overline{{\mathbf{w}}}^{\top}+\sum_{i=1}^{n}{\mathbf{y}}_{i}c_{i}^{L-1}\left({\mathbf{a}}_{L-1}^{p}\right)^{\top}.$ (115)
Hence, $\overline{{\mathbf{w}}}$ is parallel to ${\mathbf{a}}_{L-1}^{p}$. Since ${\mathbf{a}}_{L-1}$ has non-negative entries, we get that the entries of $\overline{{\mathbf{w}}}$ have the same sign. Therefore, from $|\overline{{\mathbf{w}}}|={\mathbf{a}}_{L-1}/(p\hat{p})^{1/4}$, we have $\overline{{\mathbf{w}}}=q{\mathbf{a}}_{L-1}/(p\hat{p})^{1/4}$, for some $q\in\{-1,1\}$. Also, ${\mathbf{a}}_{L-1}$ is parallel to ${\mathbf{a}}_{L-1}^{p}$. Since $p\geq 2$, and ${\mathbf{a}}_{L-1}$ has non-negative entries, the non-zero entries of ${\mathbf{a}}_{L-1}$ are identical.
Next, from the KKT conditions, we have